A 30 year mortgage bears annual interest at 12% and has a principal of $145,654. What will the monthly payments equal assuming monthly compounding
Answer #1

Principal or PV =145654
Rate per month =12%/12 =1%
Number of month =30*12 =360
Monthly Payment =PV/((1-(1+r)^-n)/r) =145654/((1-(1+1%)^-360)/1%) =1498.22